draw the line which will represent the graph of the equation y-2.5x=c if it's known it passes through the point K(2, -3)
Answer by ikleyn(52781)

.

They want you determine the value of the constant "c" in the right side of the equation.


For it, simply substitute the given coordinates 2 for x and -3 for y into the equation

    -3 - 2.5*2 = c.


Then you get c = -3 - 5 = -8.


So, your equation is  y - 2.5x = -8.    ANSWER

Solved and explained.

-------------

To make a plot, go to website www.desmos.com/calculator

Find there free of charge plotting tool for common use.

Print the line equation in the port window and get your line / plot instantly.